1. Your boss has told you to evaluate the cash flows for Machine 1 and Machine 2, using the payback period method. Which machine are you going to recommend? Justify your answer. YearlMachine 1Machine 2 ) (52,550) (40,000) 10,000 9,000 2 10,1009,025 3 10,450 9,030 4 22,000 9,045 5 20,000 9,099
Answer #1

Payback period

Year Machine 1 Commulative cash inflow Machine 2 Commulative cash inflow
0 -52550 -40000
1 10000 10000 9000 9000
2 10100 20100 9025 18025
3 10450 30550 9030 27055
4 22000 52550 9045 36100
5 20000 72550 9099 45199

Machine 1 is able to cover entire cash outflow in 4th year, hence, payback period for Machine 1 is 4 years

Machine 2, It is clearly visible from the commulative cash inflows that, Machine 2 takes slightly more than 4 years to cover its cash outflow.

Now, Payback period of Machine 2

Cash flow for 4 years = 36,100

Balance outlay = 40,000 - 36,100 = 3,900

Therefore payback period = 4 years + 3,900 / 9,099 = 4.43

Since, payback period of Machine 1 is less than Machine 2, Machine 1 should be recommended